OpenStudy (anonymous):

I'm stuck can someone help me? :c Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. y-intercept: (0,-2) m= 2 y-y1=m(x-x1) y--2=2(x-0) y+2=2x-0

OpenStudy (anonymous):

y-y1=m(x-x1) y--2=2(x-0) y+2=2x-0 thats in point slope

OpenStudy (anonymous):

i think 75% sure

OpenStudy (anonymous):

That's wrong though, it's asking me foor the line that intersects y-intercepts from the two points...

hero (hero):

@TheHero, you plug one of given points in either (3,4) or (-3,8) to the point slope formula if you want point-slope form. If I were you, I'd plug in the first point (3,4): y - y1 = m(x - x1) y - 4 = 2(x - 3) y - 4 = 2x - 6

hero (hero):

Sorry, I put the points in backwards.

OpenStudy (anonymous):

@Hero That's wrong though, it's asking me foor the line that intersects y-intercepts from the two points...

hero (hero):

The original question says to use point-slope form. Now you're bringing up something completely different.

OpenStudy (anonymous):

Write the EQUATION OF THE LINE THAT /PASSES THROUGH/ THE POINTS (3, 4) and (–3, –8) in point-slope form.

hero (hero):

Yes, and the work I wrote above answers your question.

OpenStudy (anonymous):

Maybe, I'm the one that doesn't understand it.... D: Can you explain the question to me?

hero (hero):

The point slope formula is y - y1 = m(x - x1) You plug the given point (3,4) into the formula. Then you simplify it

OpenStudy (anonymous):

Uhm ,alright then thank you! D: